# Improved real-space parallelizable matrix-product state compression and its application to unitary quantum dynamics simulation

### v1

@@ Abstract
Towards the efficient simulation of near-term quantum devices using tensor network states, we introduce an improved real-space parallelizable matrix-product state (MPS) compression method. This method enables efficient compression of all virtual bonds in constant time, irrespective of the system size, with controlled accuracy, while it maintains the stability of the wavefunction norm without necessitating sequential renormalization procedures. In addition, we introduce a parallel regauging technique to partially restore the deviated canonical form, thereby improving the accuracy of the simulation in subsequent steps. We further apply this method to simulate unitary quantum dynamics and introduce a parallel time-evolving block-decimation (pTEBD) algorithm. We employ the pTEBD algorithm for extensive simulations of typical one- and two-dimensional quantum circuits, involving over 1000 qubits. The obtained numerical results unequivocally demonstrate that the pTEBD algorithm achieves the same level of simulation precision as the current state-of-the-art MPS algorithm but in polynomially shorter time, exhibiting nearly perfect weak scaling performance on a modern supercomputer.
@@ I Introduction
Owing to the recent rapid advancement of synthetic quantum devices [1, 2, 3, 4], quantum computing is becoming a competitive candidate for next-generation computing. It utilizes these controllable quantum systems for quantum information processing and to address challenging problems beyond the capability of classical computing [5]. Due to the principles of quantum mechanics, a quantum computer maintaining a quantum state composed of $N$ qubits can access a total of $2^{N}$ dimensions in its working space, i.e., the Hilbert space, hence providing ample computational resources for addressing problems with comparable exponential complexity, such as the quantum many-body problem [6]. An ideal quantum computer can encompass the entire Hilbert space, thereby offering indisputable exponential resources compared to a classical computer, which thus establishes the quantum advantage. However, currently available quantum computers are hindered by strong decoherence noise, allowing only a small fraction of the Hilbert space to be explored [7]. These noisy quantum computers are referred to as a noisy intermediate-scale quantum (NISQ) [8] device. The central challenge in near-term development of quantum computing is determining efficient strategies for utilizing NISQ devices to achieve a practical quantum advantage in meaningful computational tasks [6, 9].
In parallel, to validate the attainment of practical quantum advantage and to establish a reliable and readily accessible testing environment for quantum algorithm development, extensive research has been focused on developing classical simulations of quantum computing. For exact simulations, two primary methods are employed: the statevector method and the tensor contraction method. The statevector method can simulate quantum circuits with any circuit depth (in polynomial time with respect to the circuit depth) and provides the complete amplitudes of the wavefunction. However, simulating a quantum circuit with more than 50 qubits using currently available classical computers is considered impossible [10, 11]. On the other hand, the tensor contraction method [12, 13, 14] can simulate quantum circuits with up to around 100 qubits, but the arrangement of gates and the circuit depth are severely restricted, allowing only a limited portion of the wavefunction amplitudes to be obtained [15]. Therefore, to efficiently simulate current NISQ devices with over 400 qubits [16], it is essential to develop more specialized simulation algorithms. These algorithms should consider the limitation that NISQ devices can only generate limited quantum entanglement.
@@
In general, an MPS-based algorithm for quantum circuit simulations includes two main procedures. One involves applying quantum gates, and the other involves compressing the MPS to a computable size. Regarding the former, we can parallelly apply multiple gates as long as they do not have spacetime overlap. For the later, it has been realized that even if all the MPS virtual bonds are compressed in parallel, the error induced by this parallel compression is still manageable [23]. These observations imply the feasibility of developing a real-space parallelizable algorithm to simulate quantum circuits. In fact, in the study of the time evolution of a quantum many-body system, a similar parallel scheme has been proposed and has obtained promising results [24]. However, a crucial issue that remains in this scheme is the fast-decaying wavefunction norm caused by parallel MPS compression, which might result in serious numerical instability. This instability necessitates the renormalization of the the wavefunction after simulating a certain period of time evolution [24]. Being a sequential procedure, this additional wavefunction renormalization significantly diminishes parallelism and hampers the efficient utilization of the parallel computing environment.
@@
Here, we introduce an improved real-space parallelizable MPS compression (IPMC) method that can stabilize the wavefunction norm without compromising parallelism. For an MPS in the canonical form, we prove that the wavefunction norm, after undergoing the IPMC, stabilizes within an interval with two bounds uniformly converging to 1 from both sides, and the convergence improves as the compression error decreases. Moreover, we integrate an additional parallel regauging procedure into the IPMC to generate a better starting point for the next simulation step, thereby enhancing the simulation accuracy. Based on this IPMC method, we propose the parallel time-evolving block-decimation (pTEBD) algorithm to simulate unitary quantum dynamics in a fully parallelizable manner. We benchmark the pTEBD algorithm by extensively simulating typical random and parametrized quantum circuits on both one-dimensional (1D) and two-dimensional (2D) qubit arrays. These numerical results demonstrate that the pTEBD algorithm achieves a comparable simulation precision to the previous sequential MPS algorithm proposed in [7] and can attain the same accuracy in polynominally shorter time. Meanwhile, the wavefunction norm stabilizes to approach 1, instead of decaying exponentially to zero, throughout the entire simulation period with a circuit depth exceeding 1000. Consequently, we successfully achieve nearly perfect weak scaling performance in the pTEBD simulation, involving over 1000 qubits.
@@
The rest of the paper is organized as follows. First, we briefly summarize the MPS representation of quantum many-body states, the properties of the canonical form, and the time-evolving block-decimation (TEBD) algorithm (also denoted as the simple update algorithm) in Sec. II. Next, we systematically describe the real-space parallelizable MPS compression method in Sec. III, where we introduce a wavefunction norm stabilization method with a proof of its bounding theorem, explore the parallel regauging technique through the trivial simple update, and finally explain the IPMC method. In Sec. IV, we propose the pTEBD algorithm and demonstrate its accuracy, numerical stability, and performance through extensive simulations of random and parametrized quantum circuits in one and two dimensions. Finally, in Sec. V, we summarize the results of this paper and briefly discuss their impact on the development of near-term quantum computing. The convergence of an MPS to its canonical form using the parallel trivial simple update is examined in Appendix A. The details of the quantum circuits used for the simulations are described in Appendix B. An additional result regarding the accuracy of the pTEBD algorithm is provided in Appendix C.
@@ II.1 Matrix-product state
where $M^{[i]\sigma_{i}}$ is a rank-$3$ tensor on site $i$ and it is simply a matrix when the index $\sigma_{i}$ is fixed. The row and column of the matrix $M^{[i]\sigma_{i}}$ are associated with the virtual space $\mathcal{V}_{i-1}$ and $\mathcal{V}_{i}$, respectively, emerged from the decomposition. The dimension of $\mathcal{V}_{i}$, $\chi_{i}=\mathrm{dim}(\mathcal{V}_{i})$, is the bond dimension of bond $i$ connecting sites $i$ and $i+1$. Especially, bond $N$ (also labeled as bond 00) connects sites $N$ and $1$. Without losing generality, we can always assign $\chi_{0}=\chi_{N}=1$ in the decomposition [19]. The corresponding MPS is usually named an open MPS. In the rest of the paper, we will only consider the open MPS.
@@ II.2 Canonical form
and the entanglement spectrum level $\xi_{\alpha_{i}}$ is given by $\xi_{\alpha_{i}}=-\ln\lambda_{\alpha_{i}}^{2}$. Therefore, truncating the bond dimension $\chi_{i}$ at bond $i$ using $\{\lambda_{\alpha_{i}}\}$ allows for the retention of a significant portion of the entanglement information between these two subsystems. For practical construction of the canonical form for a general quantum state and additional insights into its useful properties, one can find more comprehensive reviews [19, 17, 18].
@@ II.4 Computational complexities of MPS calculations
Before concluding this section, let us discuss the spacial and time complexities of generic MPS calculations. For an MPS with $N$ sites and a bond dimension $\chi$, the spacial complexity, i.e., the memory footprint, is $\mathcal{O}(N\chi^{2})$, increasing polynomially with both $N$ and $\chi$. Here, for simplicity, we assume the same bond dimension for each matrix and maintain this convention throughout the rest of the paper. Typically, the time complexity of local MPS operations in MPS-based algorithms is $\mathcal{O}(\chi^{n})$, where, for instance, $n=3$ in the density-matrix renormalization group (DMRG) algorithm [33, 34]. In addition, the time cost of real-space sequential algorithms for simulating a system with $N$ sites linearly increases with $N$, as the sweep procedure is usually required in these algorithms [19, 7]. Therefore, the complexities of MPS calculations are determined by the bond dimension $\chi$ and the system size $N$ for real-space sequential algorithms. These complexity analyses suggest that compressing an MPS to another MPS with a smaller $\chi$ while maintaining high fidelity is essential for accelerating MPS calculations.
@@ III.1 MPS compression
There are two typical approaches to performing MPS compression. One involves locally and sequentially modifying each of the MPS tensors (sequential MPS compression). Representative algorithms in this approach include the sequential SVD truncation and the sequential local variational optimization (for details, see Ref. [19]). Due to the sequential nature of this approach, the time complexity of these algorithms always linearly increases with the system size $N$. The other approach involves compressing all MPS tensors simultaneously (parallel MPS compression). Although this real-space parallelizable approach has a promising constant time complexity with increasing the system size $N$, it has been rarely discussed in the literature [23, 24, 35].
@@ III.2 Fundamental theorem on parallel MPS compression
The bound in Eq. (15) has the same formulation as the exact fidelity lower bound for the sequential SVD truncation formulated in Ref. [7]. Although $\epsilon_{i}(\chi_{i})$ may slightly deviates from the corresponding local truncation error in the sequential SVD compression, we argue that the parallel MPS compression algorithm could have the same order of accuracy as the sequential SVD compression algorithm. This argument will be numerically confirmed in Sec. IV.2.1.
@@ III.3 Wavefunction norm stabilization in parallel MPS compression
Although the norm of the wavefunction does not impact the calculation of any physical observables, it strongly influences numerical stability in actual numerical simulations, as it may diverge during the calculation process. Hence, we have to carefully consider the possible norm deviation induced by the parallel MPS compression.
@@ III.3.1 Decay of the wavefunction norm after parallel MPS compression
$n^{2}=$ $\mathrm{Tr}\{\mathcal{E}^{[1]}[P_{1}\cdots\mathcal{E}^{[N-2]}[P_{N-2}$ (20)
$\mathcal{E}^{[N-1]}(P_{N-1}\rho^{[N]}P_{N-1})P_{N-2}]\cdots P_{1}]\}$
@@
Lemma 1 shows that the wavefunction norm necessarily decreases, except for the ideal case, after a parallel MPS compression process. This implies that, if we alternatively perform TEBD procedures and parallel MPS compression processes, as in the case of Ref. [24], the wavefunction norm will monotonically decay, ultimately leading to serious numerical instability. This norm divergence will be numerically demonstrated later in Sec. IV.2.2. To avoid the numerical instability, natively, we can renormalize the wavefunction after a parallel MPS compression process. However, the time cost of calculating the wavefunction norm increases linearly with the system size $N$, offsetting the benefits of parallel MPS compression. Therefore, it is highly desired to develop an efficient method that can suppress the norm divergence during repeated parallel MPS compression processes without any sequential procedure.
@@ III.3.2 Parallelizable wavefunction norm stabilization
[…]
@@
Figure 2: Illustration of the pTEBD algorithm for a quantum circuit simulation. Starting from an MPS in (a), which is allowed to be slightly deviated from the canonical form, the state is evolved using the standard TEBD procedure but with all gates in the same layer treated parallelly in (b), and then truncate the state using the IPMC (see Sec. III.5) in (c) to update the MPS, which is subsequently used to simulate the quantum circuit with the next layer of gates. A possible realization of the distribution of dates to each MPI process ${\rm P}_{l}$ and the data transfer among MPI processes in the pTEBD algorithm is schematically shown in (d). As multiple PtSU steps may be executed in the IPMC, the data transfer is consequently required several times, as indicated here simply by two arrows pointing forward and backward between neighboring MPI processes. Note also that single-qubit gates can be treated exactly without increasing the bond dimension by any MPS-based simulations, including the pTEBD algorithm.
@@ III.4 Parallel recovery of the canonical form through parallel tSU
Because of its many advantages, almost all MPS-based algorithms developed previously employ the canonical form. Therefore, it is important to maintain an MPS in the canonical form or close to the canonical form also in a parallel MPS algorithm. In this respect, it is well-known that the right-left sweeps of tSU, which is a sequential procedure, can converge an arbitrary MPS to its canonical form. However, how the canonical form is recovered with a parallel procedure has been rarely investigated thus far. In this section, we first prove that a sufficient number of parallel tSU steps can converge an arbitrary MPS to its canonical form, and then we discuss how it approaches the canonical form with fewer operations.
@@
Then, an MPS composed of $\{\Gamma^{[i]\sigma_{i}},\Lambda^{[j]}\}$ is in the canonical form, satisfying the conditions in Eq. (II.2), is equivalent to saying that $L^{[i]\sigma_{1}\cdots\sigma_{i}}$ and $R^{[i]\sigma_{i+1}\cdots\sigma_{N}}$ satisfy
@@
Similarly, we can prove that the same PtSU steps also establish Eq. (34) for $\forall i\in[1,N-1]$. For $i=0$, Eq. (34) is automatically satisfied because of the wavefunction normalization condition. Finally, since this proof is inductive, it provides an upper bond on how many PtSU steps are required to restore the canonical form. ∎
Theorem 3 shows that the PtSU, a real-space parallelizable algorithm, can recover the canonical form with only locally operating tensors, implying the potential of its utilization on modern supercomputers where communication between neighboring nodes is highly optimized [37]. We further explore the behavior of approaching the canonical form by applying the PtSU steps in Appendix A. We numerically demonstrate that an MPS approaches the canonical form monotonically when the starting MPS is obtained from the parallel MPS compression on a canonicalized MPS. Therefore, we can append a few PtSU steps to partially recover the canonical form after a parallel MPS compression process, which improves the accuracy of subsequent TEBD evolution and MPS compression.

### v2

@@ Abstract
Towards the efficient simulation of near-term quantum devices using tensor network states, we introduce an improved real-space parallelizable matrix-product state (MPS) compression method. This method enables efficient compression of all virtual bonds in constant time, irrespective of the system size, with controlled accuracy, while it maintains the stability of the wavefunction norm without necessitating sequential renormalization procedures. In addition, we introduce a parallel regauging technique to partially restore the deviated canonical form, thereby improving the accuracy of the simulation in subsequent steps. We further apply this method to simulate unitary quantum dynamics and introduce an improved parallel time-evolving block-decimation (pTEBD) algorithm. We employ the improved pTEBD algorithm for extensive simulations of typical one- and two-dimensional quantum circuits, involving over 1000 qubits. The obtained numerical results unequivocally demonstrate that the improved pTEBD algorithm achieves the same level of simulation precision as the current state-of-the-art MPS algorithm but in polynomially shorter time, exhibiting nearly perfect weak scaling performance on a modern supercomputer.
@@ I Introduction
Owing to the recent rapid advancement of synthetic quantum devices [1, 2, 3, 4], quantum computing is becoming a competitive candidate for next-generation computing. It utilizes these controllable quantum systems for quantum information processing and to address challenging problems beyond the capability of classical computing [5]. Due to the principles of quantum mechanics, a quantum computer maintaining a quantum state composed of $N$ qubits can access a total of $2^{N}$ dimensions in its working space, i.e., the Hilbert space, hence providing ample computational resources for addressing problems with comparable exponential complexity, such as the quantum many-body problem [6]. An ideal quantum computer can encompass the entire Hilbert space, thereby offering indisputable exponential resources compared to a classical computer, which thus establishes the quantum advantage. However, currently available quantum computers are hindered by strong decoherence noise, allowing only a small fraction of the Hilbert space to be explored [7]. These noisy quantum computers are referred to as a noisy intermediate-scale quantum (NISQ) [8] device. The central challenge in near-term development of quantum computing is determining efficient strategies for utilizing NISQ devices to achieve a practical quantum advantage in meaningful computational tasks [6, 9, 10, 11].
In parallel, to validate the attainment of practical quantum advantage and to establish a reliable and readily accessible testing environment for quantum algorithm development, extensive research has been focused on developing classical simulations of quantum computing. For exact simulations, two primary methods are employed: the statevector method and the tensor contraction method. The statevector method can simulate quantum circuits with any circuit depth (in polynomial time with respect to the circuit depth) and provides the complete amplitudes of the wavefunction. However, simulating a quantum circuit with more than 50 qubits using currently available classical computers is considered impossible [12, 13]. On the other hand, the tensor contraction method [14, 15, 16, 17, 18] can simulate quantum circuits with up to around 100 qubits, but the arrangement of gates and the circuit depth are severely restricted, allowing only a limited portion of the wavefunction amplitudes to be obtained [19]. Therefore, to efficiently simulate current NISQ devices with over 400 qubits [20], it is essential to develop more specialized simulation algorithms. These algorithms should consider the limitation that NISQ devices can only generate limited quantum entanglement.
@@
In general, an MPS-based algorithm for quantum circuit simulations includes two main procedures. One involves applying quantum gates, and the other involves compressing the MPS to a computable size. Regarding the former, we can parallelly apply multiple gates as long as they do not have spacetime overlap. For the later, it has been realized that even if all the MPS virtual bonds are compressed in parallel, the error induced by this parallel compression is still manageable [27]. These observations imply the feasibility of developing a real-space parallelizable algorithm to simulate quantum circuits. In fact, in the study of the time evolution of a quantum many-body system, a similar parallel scheme has been proposed and has obtained promising results [28] (see also Ref. [29] for an alternative parallel approach to simulating Hamiltonian dynamics and Ref. [30] for a parallel algorithm of searching the ground state). However, a crucial issue that remains in this scheme is the fast-decaying wavefunction norm caused by parallel MPS compression, which might result in serious numerical instability. This instability necessitates the renormalization of the the wavefunction after simulating a certain period of time evolution [28]. Being a sequential procedure, this additional wavefunction renormalization significantly diminishes parallelism and hampers the efficient utilization of the parallel computing environment.
@@
Here, we introduce an improved real-space parallelizable MPS compression (IPMC) method that can stabilize the wavefunction norm without compromising parallelism. For an MPS in the canonical form, we prove that the wavefunction norm, after undergoing the IPMC, stabilizes within an interval with two bounds uniformly converging to 1 from both sides, and the convergence improves as the compression error decreases. Moreover, we integrate an additional parallel regauging procedure into the IPMC to generate a better starting point for the next simulation step, thereby enhancing the simulation accuracy. Based on this IPMC method, we propose the parallel time-evolving block-decimation (pTEBD) algorithm to simulate unitary quantum dynamics in a fully parallelizable manner. We benchmark the pTEBD algorithm by extensively simulating typical random and parametrized quantum circuits on both one-dimensional (1D) and two-dimensional (2D) qubit arrays. These numerical results demonstrate that the pTEBD algorithm achieves a comparable simulation precision to the previous sequential MPS algorithm proposed in [7] and can attain the same accuracy in polynominally shorter time. Meanwhile, the wavefunction norm stabilizes to approach 1, instead of decaying exponentially to zero, throughout the entire simulation period with a circuit depth exceeding 1000. Consequently, we successfully achieve nearly perfect weak scaling performance in the pTEBD simulation, involving over 1000 qubits with over 250 computational nodes on the Supercomputer Fugaku installed at RIKEN.
@@
The rest of the paper is organized as follows. First, we briefly summarize the MPS representation of quantum many-body states, the properties of the canonical form, and the time-evolving block-decimation (TEBD) algorithm (also denoted as the simple update algorithm) in Sec. II. Next, we systematically describe the real-space parallelizable MPS compression method in Sec. III, where we introduce a wavefunction norm stabilization method with a proof of its bounding theorem, explore the parallel regauging technique through the trivial simple update, and finally explain the IPMC method. In Sec. IV, we propose the pTEBD algorithm and demonstrate its accuracy, numerical stability, and performance through extensive simulations of random and parametrized quantum circuits in one and two dimensions. Finally, in Sec. V, we summarize the results of this paper and briefly discuss their impact on the development of near-term quantum computing. The details of the quantum circuits used for the simulations are described in Appendix A. An additional result regarding the accuracy of the pTEBD algorithm is provided in Appendix B. Furthermore, the pTEBD algorithm is applied to the simulation of the quantum Fourier transformation in Appendix C.
@@ II.1 Matrix-product state
where $M^{[i]\sigma_{i}}$ is a rank-$3$ tensor on site $i$ and it is simply a matrix when the index $\sigma_{i}$ is fixed. The row and column of the matrix $M^{[i]\sigma_{i}}$ are associated with the virtual spaces $\mathcal{V}_{i-1}$ and $\mathcal{V}_{i}$, respectively, emerged from the decomposition. The dimension of $\mathcal{V}_{i}$, $\chi_{i}=\mathrm{dim}(\mathcal{V}_{i})$, is the bond dimension of bond $i$ connecting sites $i$ and $i+1$. Especially, bond $N$ (also labeled as bond 00) connects sites $N$ and $1$. Without losing generality, we can always assign $\chi_{0}=\chi_{N}=1$ in the decomposition [23]. The corresponding MPS is usually named an open MPS. In the rest of the paper, we will only consider the open MPS.
@@ II.2 Canonical form
and the entanglement spectrum level $\xi_{\alpha_{i}}$ is given by $\xi_{\alpha_{i}}=-\ln\lambda_{\alpha_{i}}^{2}$. Therefore, truncating the bond dimension $\chi_{i}$ at bond $i$ on the basis of $\{\lambda_{\alpha_{i}}\}$ allows for the retention of a significant portion of the entanglement information between these two subsystems. For practical construction of the canonical form for a general quantum state and additional insights into its useful properties, one can find more comprehensive reviews [23, 21, 22].
@@ II.4 Computational complexities of MPS calculations
Before concluding this section, let us discuss the spacial and time complexities of generic MPS calculations. For an MPS with $N$ sites and a bond dimension $\chi$, the spacial complexity, i.e., the memory footprint, is $\mathcal{O}(N\chi^{2})$, increasing polynomially with both $N$ and $\chi$. Here, for simplicity, we assume the same bond dimension for each matrix and maintain this convention throughout the rest of the paper. Typically, the time complexity of local MPS operations in MPS-based algorithms is $\mathcal{O}(\chi^{n})$, where, for instance, $n=3$ in the TEBD algorithm and the density-matrix renormalization group (DMRG) algorithm [39, 40]. In addition, the time cost of real-space sequential algorithms for simulating a system with $N$ sites linearly increases with $N$, as the sweep procedure is usually required in these algorithms [23, 7]. Therefore, the complexities of MPS calculations are determined by the bond dimension $\chi$ and the system size $N$ for real-space sequential algorithms. These complexity analyses suggest that compressing an MPS to another MPS with a smaller $\chi$ while maintaining high fidelity is essential for accelerating MPS calculations. Notice that we omit the dimension $d_{\rm loc}$ of the local Hilbert space in the complexity analysis since it is in general small (in our case, $d_{\rm loc}=2$) compared with $\chi$ and $N$.
@@ III.1 MPS compression
There are two typical approaches to performing MPS compression. One involves locally and sequentially modifying each of the MPS tensors (sequential MPS compression). Representative algorithms in this approach include the sequential SVD truncation and the sequential local variational optimization (for details, see Ref. [23]). In this case, $\mathcal{F}_{2}$ in Eq. (12) can be evaluated from the singular values kept in $|\Psi_{T}\rangle$. Due to the sequential nature of this approach, the time complexity of these algorithms always linearly increases with the system size $N$. The other approach involves compressing all MPS tensors simultaneously (parallel MPS compression). Although this real-space parallelizable approach has a promising constant time complexity with increasing the system size $N$, it has been rarely discussed in the literature [27, 28, 41]. In this parallel approach, it is not straightforward to build the connection between the fidelities in each local tensor compression and the global fidelity in Eq. (12).
@@ III.2 Fundamental theorem on parallel MPS compression
Note that the right-hand side of Eq. (15) can be negative, while the wavefunction fidelity $\mathcal{F}_{2}$ must be positive. The bound in Eq. (15) has the same formulation as the exact fidelity lower bound for the sequential SVD truncation formulated in Ref. [7]. Although $\epsilon_{i}(\chi_{i})$ may slightly deviates from the corresponding local truncation error in the sequential SVD compression, we argue that the parallel MPS compression algorithm could have the same order of accuracy as the sequential SVD compression algorithm. This argument will be numerically confirmed in Sec. IV.2.1. Also notice that, in general, the canonical form is broken after a parallel MPS compression, which could impact on the subsequent simulation steps.
@@ III.3 Wavefunction norm stabilization in parallel MPS compression
Although the norm of the wavefunction does not impact the calculation of any physical observables, it strongly influences numerical stability in actual numerical simulations, as it may vanish during the calculation process. Hence, we have to carefully consider the possible norm deviation induced by the parallel MPS compression.
@@ III.3.1 Decay of the wavefunction norm after parallel MPS compression
$n^{2}=$ $\mathrm{Tr}[\mathcal{E}^{[1]}(P_{1}\cdots\mathcal{E}^{[N-2]}(P_{N-2}$ (20)
$\mathcal{E}^{[N-1]}(P_{N-1}\rho^{[N]}P_{N-1})P_{N-2})\cdots P_{1})]$
@@
Lemma 1 shows that the wavefunction norm necessarily decreases, except for the ideal case, i.e., without truncating any bonds, after a parallel MPS compression process. This implies that, if we alternatively perform TEBD procedures and parallel MPS compression processes, as in the case of Ref. [28], the wavefunction norm will monotonically decay, ultimately leading to serious numerical instability. This norm vanishing will be numerically demonstrated later in Sec. IV.2.2. To avoid the numerical instability, naively, we can renormalize the wavefunction after a parallel MPS compression process. However, the time cost of calculating the wavefunction norm increases linearly with the system size $N$, offsetting the benefits of parallel MPS compression. Therefore, it is highly desired to develop an efficient method that can suppress the norm vanishing during repeated parallel MPS compression processes without any sequential procedure.
@@ III.3.2 Parallelizable wavefunction norm stabilization
$n^{*}_{\text{l}ower}$ $=$ $(1-\sqrt{2\epsilon(\chi^{\prime})})\prod_{i=1}^{N-1}\nu_{i}~{},$ (27)
$n^{*}_{\text{u}pper}$ $=$ $\prod_{i=1}^{N-1}\nu_{i}~{},$ (28)
with $n^{*}_{\text{l}ower}\leq 1$ and $n^{*}_{\text{u}pper}\geq 1$. Moreover, they uniformly converge to 1 as $\epsilon(\chi^{\prime})\rightarrow 0$.
@@ III.4 Parallel recovery of the canonical form through parallel tSU
Because of its many advantages, almost all MPS-based algorithms developed previously employ the canonical form. Therefore, it is important to maintain an MPS in the canonical form or close to the canonical form also in a parallel MPS algorithm. In this respect, it is well-known that the right-left sweeps of tSU, which is a sequential procedure, can converge an arbitrary MPS to its canonical form. However, how the canonical form is recovered with a parallel procedure has been rarely investigated thus far. In this section, we first prove that at most $\frac{N}{2}$ ($\frac{N-1}{2}$) parallel tSU steps can converge an arbitrary MPS to its canonical form when $N$ is even (odd), and then we discuss how it approaches the canonical form with fewer operations.
@@
Then, asserting that an MPS composed of $\{\Gamma^{[i]\sigma_{i}},\Lambda^{[j]}\}$ is in the canonical form, satisfying the conditions in Eq. (II.2), is equivalent to saying that $L^{[i]\sigma_{1}\cdots\sigma_{i}}$ and $R^{[i]\sigma_{i+1}\cdots\sigma_{N}}$ satisfy
@@
Similarly, we can prove that the same PtSU steps also establish Eq. (34) for $\forall i\in[1,N-1]$. For $i=0$, Eq. (34) is automatically satisfied because of the wavefunction normalization condition. Finally, since this proof is inductive, it provides an upper bound on how many PtSU steps are required to restore the canonical form. ∎
Theorem 3 shows that the PtSU, a real-space parallelizable algorithm, can recover the canonical form with only locally operating tensors, implying the potential of its utilization on modern supercomputers where communication between neighboring nodes is highly optimized [43]. Moreover, we can simply skip applying the identity gate in the tSU procedure to further optimize the performance in practice.
Now, we explore the behavior of approaching the canonical form by applying the PtSU. To measure how an MPS approaches its canonical form with PtSU steps, we first define the following distance to the canonical form:
where |||||$\cdot$||| | ⋅ | | represents the Frobenius norm of the matrix. This quantity is zero if and only if an MPS is in the canonical form. Note that, in Eq. (36), we assume that the norm of the MPS is 1. Notice that a similar quantity has been proposed in a recent study to examine the regauging of general tensor networks, where the nuclear norm is considered [44].
We investigate how $\mathcal{C}$ changes with the application of PtSU steps. To closely mimic the actual calculation task, we prepare the initial MPS as follows: We first construct a canonicalized random MPS with the bond dimension $\chi$. Each tensor element in this MPS is initialized to a complex number $z=a+ib$ with $a$ and $b$ chosen randomly from $[-1,1]$, and then its canonical form is constructed with the norm properly normalized to 1. Subsequently, we truncate all the virtual bonds to $\chi/2$, i.e., disregarding the $\chi/2$ smallest elements in each $\Lambda^{[i]}$, and renormalize the wavefunction norm to 1. We then perform PtSU steps on this MPS and calculate $\mathcal{C}$ after each PtSU step. The results are summarized in Fig. 2. Here, we choose $N=20$ and 24 and examine the cases of $\chi=32$, 128, 256 and 512.
Figure 2: Seim-log plot of $\mathcal{C}$ as a function of PtSU steps for $N=20$ and 24. The initial MPS is prepared by simultaneously truncating all the virtual bonds of a canonicalized random MPS to half of the bond dimension and manually normalizing the norm of the resulting MPS to 1. The results are averaged over 100 initial MPSs with different sets of random parameters.
In Fig. 2, we observe that $\mathcal{C}$ rapidly and monotonically decreases with PtSU steps, especially at the beginning of several PtSU steps. Additionally, the canonical form is attained more rapidly when the bond dimension is larger. This implies that we can drive the MPS closer to the canonical form by appending a few PtSU steps after the parallel MPS compression. Also notice that the canonical form is exactly restored after $N/2$ PtSU steps, justifying Theorem 3.